Express 2x+1 / x(x+1) in partial fractions and hence find the general solution of the differential equation x(x+1)dy/dx = y(2x+1) expressing y explicitly in terms of x.


1
Expert's answer
2021-09-15T02:57:08-0400

It follows that "\frac{2x+1}{x(x+1)}=\frac{1}{x}+\frac{1}{x+1}" is the expression in partial fractions. Let us find the general solution of the differential equation "x(x+1)\frac{dy}{dx} = y(2x+1)" expressing "y" explicitly in terms of "x". We get that "\frac{dy}{y} = \frac{2x+1}{x(x+1)}dx," and hence "\int\frac{dy}{y} = \int\frac{2x+1}{x(x+1)}dx=\int(\frac{1}{x}+\frac{1}{x+1})dx." We conclude that "\ln|y|=\ln|x|+\ln|x+1|+\ln|C|=\ln|x(x+1)C|." Therefore, "y=Cx(x+1)" is the general solution of the differential equation "x(x+1)\frac{dy}{dx} = y(2x+1)".